# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

### v2

@@ Appendix A Proof of Lemma 6
where the right-hand side gives the tight $\delta(\varepsilon)$ for the Gaussian mechanism with sensitivity $\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ and noise scale $\sigma$. Further, since $V_{r}$ has orthonormal columns,
@@
$=\left\lVert A^{+}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$
@@
Appendix B Proof of Lemma 7
@@
$=H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right).$
By the compact SVD $H_{T}=U_{r}\Sigma_{r}V_{r}^{\top}$, where $r$ denotes the rank of $H_{T}$, we have that for all $\alpha\geq 0$,
$H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
@@
$=$ $H_{\alpha}\left(\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
@@
where $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$. ∎
Appendix C Proof of Thm. 8
Proof.
@@
$\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}=(\Delta x)^{\top}V_{r}V_{r}^{\top}(\Delta x)\leq\mathbf{1}_{T}^{\top}\left|V_{r}V_{r}^{\top}\right|\mathbf{1}_{T}.$
Appendix D Proof of Lemma 11
@@
Since $W$ is primitive and row-stochastic, $1$ is a simple eigenvalue with right eigenvector $\mathbf{1}$ and left eigenvector $\pi^{\top}$, and all other eigenvalues satisfy $|\lambda|<1$ (Perron-Frobenius). Choose a basis matrix S=[ 1B]S=[\ $\mathbf${1}\ \ B\ ] for some $B\in\mathbb{R}^{n\times n-1}$ so that the first row of $S^{-1}$ is $\pi^{\top}$. Then
$S^{-1}WS\;=\;\begin{bmatrix}1&0\\ 0&T^{\prime}\end{bmatrix}$,
and taking powers gives
$W^{t}\;=\;S\begin{bmatrix}1&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\;=\;\underbrace{\mathbf{1}\,\pi^{\top}}_{=\ \Pi}\;+\;\underbrace{S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}}_{=:R^{t}}.$
To bound $\|R^{t}\|$, fix any matrix norm $\|\cdot\|$. Since $\rho(T^{\prime}):=\max\{|\lambda|:\lambda\in\mathrm{spec}(T^{\prime})\}<1$, the Gelfand formula implies that for any $\rho$ with $\rho(T^{\prime})<\rho<1$ there exists $C<\infty$ such that $\|(T^{\prime})^{t}\|\leq C\,\rho^{t}$ for all $t\geq 1$. Therefore
$\|R^{t}\|\;=\;\left\|S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\right\|\;\leq\;\|S\|\,\|S^{-1}\|\,\|(T^{\prime})^{t}\|\;\leq\;\big(\|S\|\,\|S^{-1}\|\;C\big)\,\rho^{t}.$
Setting $M:=\|S\|\,\|S^{-1}\|\,C$ gives the claim. ∎
Appendix E Proof Theorem 13

### v3

@@ Appendix B Proof of Lemma 6
where the right-hand side gives the $\delta(\varepsilon)$-privacy profile for the Gaussian mechanism with sensitivity $\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ and noise scale $\sigma$. Further, since $V_{r}$ has orthonormal columns,
@@
$=\left\lVert A^{+}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2},$
@@
Appendix C Proof of Lemma 9
Proof.
Since $W$ is symmetric, $W=Q\Lambda Q^{\top}$ with $Q$ orthogonal and $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})$ real. Doubly stochastic implies $\lambda_{1}=1$ with eigenvector $\mathbf{1}$. Primitivity implies $|\lambda_{k}|<1$ for all $k\geq 2$. Let $P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ be the orthogonal projector onto $\mathrm{span}\{\mathbf{1}\}$ and set R:=WPR:=W-P. Then $RP=PR=0$ and $\|R^{t}\|_{2}=\max_{k\geq 2}|\lambda_{k}|^{t}$. With $\rho:=\max_{k\geq 2}|\lambda_{k}|\in(0,1)$ the claim follows. ∎
Appendix D Proof of Lemma 10
@@
$=H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right)$
which gives the claim.
Appendix E Proof of Thm. 12
Proof.
The claim is obtained by applying Lemma 6 (DP guarantees of a projected Gaussian mechanism via Moore–Penrose pseudoinverse) to the characterization given by Lemma 10. For completeness, we apply the pseudoinverse $H_{T}^{+}$ step by step. By Lemma 10, the privacy guarantees are obtained via the hockey-stick divergence $H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right)$, $\alpha\geq 0$. By the compact SVD $H_{T}=U_{r}\Sigma_{r}V_{r}^{\top}$, where $r$ denotes the rank of $H_{T}$, we have that for all $\alpha\geq 0$,
$H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right)$
@@
where
$c=\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}$
and $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$. ∎
@@
$\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}=(\Delta x)^{\top}V_{r}V_{r}^{\top}(\Delta x)\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|V_{r}V_{r}^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
where $\left|\cdot\right|$ denotes the element-wise abolute value of a matrix. ∎
Appendix F Numerical Illustration of the Spectral Estimate of Sec. 3.4
We next consider a numerical illustration of the upper bound for the sensitivity
$\Delta^{T}_{j\rightarrow\mathcal{A}}=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2},$
where $H_{T}$ is the system matrix given in Lemma 10. We compare the upper bound based on the spectral estimate $\lambda_{\max}(M_{T})$ to a lower bound for the sensitivity $\Delta^{T}_{j\rightarrow\mathcal{A}}$ we obtain with the choice $c=\mathbf{1}=\begin{bmatrix}1&\ldots&1\end{bmatrix}^{\top}$. We compare the bounds in the same two cases considered in Section 4.1: we first consider a randomly drawn synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$, and second, a network generated using a preferential-attachment model (Barabási and Albert, 1999) with $n=100$ nodes, initialized from a fully connected core of 5 nodes, where each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees. And in both cases, we take the gossip matrix $W$ to be a doubly stochastic matrix obtained from so called Metropolis–Hastings averaging of a given symmetric adjacency matrix.
As illustrated by Figure 2 the upper bound is in both cases less than 10 percent higher than the actual tight upper bound.
Figure 2: The upper bound for the squared and scaled sensitivity $(\Delta^{T}(j\to i))^{2}/T$ as a function of $T$ for a randomly drawn Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$ and b) a randomly drawn doubly-stochastic gossip matrix of a preferential-attachment model with $n=100$ nodes.
Appendix G Proof of Theorem 13
We next give proof of Theorem 13. In the proof, Lemma G.2 (Section G.2) contains a crucial step: it shows that the transient component of the sensitivity can be bounded independent of $T$.
G.1 Proof of Theorem 13
Theorem G.1.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic, and primitive. Let $S=e_{i}^{\top}$ be the single-node observation matrix. Then for any $j\in\{1,\dots,n\}$,
$\lim_{T\to\infty}\frac{\Delta_{T}(j\to i)^{2}}{T}=\frac{1}{n},$
where
$\Delta_{T}(j\to i)^{2}:=\max_{c\in\{\pm 1\}^{T}}\left\|H_{T}^{+}H_{T}(c\otimes e_{j})\right\|_{2}^{2}.$
@@ G.1 Proof of Theorem 13
We start by stating the sensitivity formula as a quadratic form. For $c\in\{\pm 1\}^{T}$ define
$v(c):=c\otimes e_{j},\qquad b(c):=H_{T}v(c)\in\mathbb{R}^{T},$
and set
$K_{T}:=H_{T}H_{T}^{\top}.$
Since $H_{T}^{+}=H_{T}^{\top}(H_{T}H_{T}^{\top})^{-1}$ and $H_{T}$ has full row rank, $K_{T}\succ 0$ and
$\|H_{T}^{+}H_{T}v(c)\|_{2}^{2}=b(c)^{\top}K_{T}^{-1}b(c),$
hence
$\Delta_{T}(j\to i)^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}K_{T}^{-1}b(c).$ (G.1)
Denoting $a_{m}:=(W^{m})_{ij}$ for $m\in\mathbb{N}$, we have for $t=1,\dots,T$,
$b_{t}(c)=\sum_{k=1}^{t}a_{t-k}c_{k}.$ (G.2)
We start by decomposing $K_{T}$. By Lemma 9, we can write
$W=P+R,\qquad P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top},\qquad\mathbf{1}^{\top}R=0,\qquad RP=PR=0,$ (G.3)
and $\|R^{t}\|_{2}\leq\rho^{t}$ for $\rho:=\max_{k\geq 2}|\lambda_{k}(W)|$. Using (G.3) together with the block-Toeplitz structure of $H_{T}$, we obtain the decomposition
$K_{T}=\frac{1}{n}J+G_{T},$ (G.4)
where $G_{T}\succeq 0$ and $J\succ 0$ satisfies $J=LL^{\top}$, with $L$ the $T\times T$ lower-triangular all-ones matrix:
$L=\begin{bmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 1&1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&1&1&\cdots&1\end{bmatrix}\in\mathbb{R}^{T\times T}.$
Equivalently, $J_{t,s}=\min(t,s)$. Therefore,
$K_{T}\succeq\frac{1}{n}J$
which further implies that
$K_{T}^{-1}\preceq nJ^{-1}.$ (G.5)
Let $D$ be the backward difference matrix on $\mathbb{R}^{T}$,
$(Dx)_{1}=x_{1},\qquad(Dx)_{t}=x_{t}-x_{t-1}\ \ (t\geq 2),$
i.e.,
$D=\begin{bmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&0&\cdots&-1&1\end{bmatrix}\in\mathbb{R}^{T\times T}.$
Since $L^{-1}=D$ and $J^{-1}=D^{\top}D$, for all $x\in\mathbb{R}^{T}$,
$x^{\top}J^{-1}x=\|Dx\|_{2}^{2}.$ (G.6)
We next split $b(c)$ into stationary and transient parts. Define the stationary and transient terms as
$a_{m}^{(P)}:=(P)_{ij}=\frac{1}{n}\quad(m\geq 1),\qquad a_{m}^{(R)}:=(R^{m})_{ij}\quad(m\geq 1),$ (G.7)
and note $a_{0}=(W^{0})_{ij}=\mathbf{1}\{i=j\}$. Accordingly, write
$b(c)=b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c),$ (G.8)
where $b^{(0)}$ denotes the (possible) $a_{0}$ contribution,
$b_{t}^{(P)}(c):=\sum_{k=1}^{t-1}\frac{1}{n}\,c_{k}=\frac{1}{n}\sum_{k=1}^{t-1}c_{k},$ (G.9)
and
$b_{t}^{(R)}(c):=\sum_{k=1}^{t-1}(R^{t-k})_{ij}c_{k}=\sum_{m=1}^{t-1}(R^{m})_{ij}\,c_{t-m}.$ (G.10)
We next derive an upper bound for the quadratic form $b(c)^{\top}K_{T}^{-1}b(c)$. For any $c$,
$b(c)^{\top}K_{T}^{-1}b(c)$ $=\|b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c)\|_{K_{T}^{-1}}^{2}$
$\leq\Big(\|b^{(0)}(c)\|_{K_{T}^{-1}}+\|b^{(P)}(c)\|_{K_{T}^{-1}}+\|b^{(R)}(c)\|_{K_{T}^{-1}}\Big)^{2}.$
The term $b^{(0)}(c)$ contributes at most an $O(1)$ additive constant, so we ignore it in the asymptotics. By (G.5) and (G.6),
$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}\leq n\,\|Db^{(P)}(c)\|_{2}^{2}.$
From (G.9), $(Db^{(P)}(c))_{t}=\frac{1}{n}c_{t-1}$ for $t\geq 2$ and $(Db^{(P)}(c))_{1}=0$, hence $\|Db^{(P)}(c)\|_{2}^{2}=(T-1)/n^{2}$ and therefore
$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}\leq\frac{T-1}{n}.$ (G.11)
By Lemma G.2 given in Section G.2, $\|b^{(R)}(c)\|_{K_{T}^{-1}}^{2}\leq C_{\mathrm{mix}}(W)$. Thus, for all $c$,
$b(c)^{\top}K_{T}^{-1}b(c)\leq\left(\sqrt{\frac{T-1}{n}}+\sqrt{C_{\mathrm{mix}}(W)}+O(1)\right)^{2}=\frac{T}{n}+O(\sqrt{T}).$
Taking the maximum over $c$ gives
$\limsup_{T\to\infty}\frac{\Delta_{T}(j\to i)^{2}}{T}\leq\frac{1}{n}.$ (G.12)
We derive a lower bound using $c=\mathbf{1}$, i.e., $c$ being all ones. Then
$(Db)_{t}=b_{t}-b_{t-1}=(W^{t-1})_{ij}=\frac{1}{n}+r_{t-1}.$
Hence
$\|Db\|_{2}^{2}=\sum_{t=1}^{T}\left(\frac{1}{n}+r_{t-1}\right)^{2}=\frac{T}{n^{2}}+\frac{2}{n}\sum_{t=1}^{T}r_{t-1}+\sum_{t=1}^{T}r_{t-1}^{2}.$
Using $|r_{t-1}|\leq\rho^{t-1}$ and summing geometric series, we have
$\left|\sum_{t=1}^{T}r_{t-1}\right|\leq\sum_{t=1}^{\infty}\rho^{t-1}=\frac{1}{1-\rho}$
and
$\sum_{t=1}^{T}r_{t-1}^{2}\leq\sum_{t=1}^{\infty}\rho^{2(t-1)}=\frac{1}{1-\rho^{2}}.$
Therefore,
$\|Db\|_{2}^{2}\geq\frac{T}{n^{2}}-\frac{2}{n(1-\rho)}-\frac{1}{1-\rho^{2}}.$ (G.13)
Next, recall that
$K_{T}=H_{T}H_{T}^{\top}=\frac{1}{n}J+G_{T},$
where $J=LL^{\top}$ with $J_{t,s}=\min(t,s)$, and $G_{T}\succeq 0$, and $J^{-1}=D^{\top}D$, where $D$ is the backward difference operator.
Since $K_{T}$ is positive definite, we use the variational representation
$b^{\top}K_{T}^{-1}b=\max_{x\in\mathbb{R}^{T}}\left(2b^{\top}x-x^{\top}K_{T}x\right).$
We lower bound this expression using
$x_{0}:=nJ^{-1}b=nD^{\top}Db.$
Then
$b^{\top}K_{T}^{-1}b$ $\geq 2b^{\top}x_{0}-x_{0}^{\top}K_{T}x_{0}$
$=2n\,b^{\top}J^{-1}b-x_{0}^{\top}\left(\tfrac{1}{n}J+G_{T}\right)x_{0}$
$=n\,b^{\top}J^{-1}b-x_{0}^{\top}G_{T}x_{0}$
$=n\|Db\|_{2}^{2}-x_{0}^{\top}G_{T}x_{0}.$
Thus, the required task is to bound the term $x_{0}^{\top}G_{T}x_{0}$. Write
$W=\frac{1}{n}11^{\top}+R$
with $1^{\top}R=0$ and $\|R^{t}\|_{2}\leq\rho^{t}$. This gives a decomposition
$H_{T}=H_{T}^{1}+H_{T}^{R},$
where $H_{T}^{1}$ is generated by $\frac{1}{n}11^{\top}$ and $H_{T}^{R}$ by $R$. Consequently,
$G_{T}=H_{T}^{R}(H_{T}^{R})^{\top}.$
Since $\|R^{t}\|_{2}\leq\rho^{t}$, a geometric-series argument gives (the 2-norm if defined by the first block column of the block lower-triangular matrix $H_{T}^{R}$)
$\|H_{T}^{R}\|_{2}\leq\frac{1}{1-\rho}.$
Recall that
$x_{0}=nD^{\top}Db,\qquad(Db)_{t}=\frac{1}{n}+r_{t-1},\qquad|r_{t}|\leq\rho^{t}.$
Let $y:=Db$. We then have that
$(D^{\top}y)_{t}=y_{t}-y_{t+1}\quad(1\leq t\leq T-1),\qquad(D^{\top}y)_{T}=y_{T}.$
Hence, for $1\leq t\leq T-1$,
$(x_{0})_{t}=n(y_{t}-y_{t+1})=n(r_{t-1}-r_{t}),$
and
$(x_{0})_{T}=ny_{T}=1+nr_{T-1}.$
Using $|r_{t-1}-r_{t}|\leq|r_{t-1}|+|r_{t}|\leq(1+\rho)\rho^{t-1}$, we obtain
$\sum_{t=1}^{T-1}(x_{0})_{t}^{2}\leq n^{2}(1+\rho)^{2}\sum_{t=1}^{\infty}\rho^{2(t-1)}=n^{2}\frac{(1+\rho)^{2}}{1-\rho^{2}}=n^{2}\frac{1+\rho}{1-\rho}.$
Moreover,
$(x_{0})_{T}^{2}=(1+nr_{T-1})^{2}\leq(1+n)^{2}.$
Combining these gives
$\|x_{0}\|_{2}^{2}\leq(1+n)^{2}+n^{2}\frac{1+\rho}{1-\rho},$
and furthermore
$x_{0}^{\top}G_{T}x_{0}\leq\frac{(1+n)^{2}}{(1-\rho)^{2}}+n^{2}\frac{1+\rho}{(1-\rho)^{3}}.$ (G.14)
Combining the bounds of Eq. (G.13) and (G.14) gives
$b^{\top}K_{T}^{-1}b\geq\frac{T}{n}-\frac{2}{n(1-\rho)}-\frac{1}{1-\rho^{2}}-\frac{(1+n)^{2}}{(1-\rho)^{2}}-n^{2}\frac{1+\rho}{(1-\rho)^{3}}.$
Since the particular choice of $c=\mathbf{1}$ gives a lower bound, we see that
$\limsup_{T\to\infty}\frac{\Delta_{T}(j\to i)^{2}}{T}\geq\frac{1}{n}.$
This together with (G.12) completes the proof. ∎
G.2 Bound for the Transient Term
We next prove a technical lemma required for the proof of Thm. 13. This lemma shows that the transient terms stay bounded independent of $T$.
Lemma G.2 (Uniform bound for the transient component).
Let $c\in\{-1,1\}^{T}$ and let $b^{(R)}(c)$ and $K_{T}$ be defined as above in the proof of Thm. 13. There exists a constant $C_{\mathrm{mix}}(W)<\infty$ depending only on $W$, such that for all $T\geq 1$ and all $c\in\{\pm 1\}^{T}$,
$b^{(R)}(c)^{\top}K_{T}^{-1}b^{(R)}(c)\leq C_{\mathrm{mix}}(W),$
where $C_{\mathrm{mix}}(W)=\frac{\pi^{2}}{6}(1-\rho)^{-4}$ and $\rho:=\max_{k\geq 2}|\lambda_{k}|$.